Norms of Composition Operators on the Hardy Space

John J. Thrall

Abstract

We show that the norm of a composition operator on the classical Hardy space H^{2} cannot be computed using only the set of H^{2} reproducing kernels, answering a question raised by Cowen and MacCluer.